OpenStudy (gabylovesyou):

Choose the correct simplification of the expression (6x2y)2(y2)3. 12x4y8 36x4y5 36x4y8 12x2y7 Question 14 (Multiple Choice Worth 5 points) [7.02] Choose the correct sum of the polynomials (6x3 − 8x − 5) + (3x3 + 6x + 2). 9x3 − 2x − 3 3x3 + 2x + 3 3x3 − 2x − 3 9x3 + 14x + 3

jimthompson5910 (jim_thompson5910):

(6x^2)^2 simplifies to what

OpenStudy (gabylovesyou):

36x^4

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

oh sry there was a y in there as well

OpenStudy (gabylovesyou):

yes

jimthompson5910 (jim_thompson5910):

(6x^2y)^2 turns into 3x^4y^2

OpenStudy (gabylovesyou):

how ?

jimthompson5910 (jim_thompson5910):

(y^2)^3 = y^6 because you multiply exponents so (6x^2y)^2*(y^2)^3 turns into 36x^4y^2*y^6 then you add the exponents for the y terms to get 2+6 = 8

jimthompson5910 (jim_thompson5910):

because y is the same as y^1

jimthompson5910 (jim_thompson5910):

then you multiply exponents

jimthompson5910 (jim_thompson5910):

afterwards 36x^4y^2*y^6 turns into 36x^4y^8

jimthompson5910 (jim_thompson5910):

do you see how I got this?

OpenStudy (gabylovesyou):

kinda.. how did u simplify(6x^2y)^2

jimthompson5910 (jim_thompson5910):

here's another way to see it

jimthompson5910 (jim_thompson5910):

(6x^2y)^2 (6x^2y)*(6x^2y) (6*6)*(x^2*x^2)*(y*y) 36x^4y^2

jimthompson5910 (jim_thompson5910):

So (6x^2y)^2 simplifies to 36x^4y^2

OpenStudy (gabylovesyou):

oh ok ! got it!

OpenStudy (gabylovesyou):

the answer is 36x4y8

OpenStudy (gabylovesyou):

thank you

jimthompson5910 (jim_thompson5910):

yes it is

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (gabylovesyou):

oh wait i forgot the other question lol

jimthompson5910 (jim_thompson5910):

6x^3 + 3x^3 = ???

OpenStudy (gabylovesyou):

9x6?

jimthompson5910 (jim_thompson5910):

no you don't add the exponents

jimthompson5910 (jim_thompson5910):

just the coefficients

OpenStudy (gabylovesyou):

ohhhh so i leave 3?

jimthompson5910 (jim_thompson5910):

-8x+6x = ???

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (gabylovesyou):

-2x

jimthompson5910 (jim_thompson5910):

-5 + 2 = ??

OpenStudy (gabylovesyou):

-3

jimthompson5910 (jim_thompson5910):

put this all together to get 9x^3 - 2x - 3

OpenStudy (gabylovesyou):

thank you!!!!

jimthompson5910 (jim_thompson5910):

you're welcome